# Tip growth in a strongly concentrated aggregation model follows local geodesics

### v2

@@ 4.1 Lipschitz bound on the locations and probabilities
Next, by the chain rule and by definition of $q_{t}(z)$,
$\frac{\mathrm{d}}{\mathrm{d}t}\Phi_{t}^{\prime\prime}(x_{j}(t))$ $=q_{t}^{\prime\prime}(x_{j}(t))+\Phi_{t}^{(3)}(x_{j}(t))\frac{\mathrm{d}}{\mathrm{d}t}x_{j}(t).$
When $\xi_{t}\not=x_{j}(t)$, using (4.3) we can calculate
$\frac{\mathrm{d}}{\mathrm{d}t}\Phi_{t}^{\prime\prime}(x_{j}(t))$ $=2\left(1-\frac{2\xi_{t}^{2}}{(x_{j}(t)-\xi_{t})^{2}}\right)\Phi_{t}^{\prime\prime}(x_{j}(t)),$
so $\frac{\frac{\mathrm{d}}{\mathrm{d}t}\Phi_{t}^{\prime\prime}(x_{j}(t))}{\Phi_{t}^{\prime\prime}(x_{j}(t))}=\exp(O(1))$. When $\xi_{t}=x_{j}(t)$, we will compute $(\Phi_{t}\circ f_{\phi^{j}_{t},c})^{\prime\prime}(x_{j}(t))$ for $0<c<\mathbf{c}$. Since $f_{\phi^{j}_{t},c}^{\prime}(x_{j}(t))=0$, the chain rule gives
$(\Phi_{t}\circ f_{\phi_{t}^{j},c})^{\prime\prime}(x_{j}(t))$ $=f_{\phi^{j}_{t},c}^{\prime\prime}(x_{j}(t))\Phi_{t}^{\prime}(f_{\phi^{j}_{t},c}(x_{j}(t))).$ (4.4)
Differentiating the expression for $f^{\prime}$ in (2.2) gives
$f_{\phi_{t}^{j},c}^{\prime\prime}(x_{j}(t))=\frac{1}{x_{j}(t)}\frac{1+d(c)}{2\sqrt{1-e^{-c}}}.$ (4.5)
As $\Phi_{t}^{\prime}(x_{j}(t))=0$ and $f_{\phi_{t}^{j},c}(x_{j}(t))-x_{j}(t)=x_{j}(t)d(c)$, we also have
$\Phi_{t}^{\prime}(f_{\phi_{t}^{j},c}(x_{j}(t)))=x_{j}(t)d(c)\Phi_{t}^{\prime\prime}(x_{j}(t))+\frac{x_{j}(t)^{2}d(c)^{2}}{2}\Phi_{t}^{(3)}(x_{j}(t))+O(d(c)^{3}).$
Since $\Phi_{t}^{(3)}(x_{j}(t))=-3\frac{\Phi_{t}^{\prime\prime}(x_{j}(t))}{x_{j}(t)}$ and $d(c)=O(c^{1/2})$, we have
$\Phi_{t}^{\prime}(f_{\phi_{t}^{j},c}(x_{j}(t)))$ $=x_{j}(t)d(c)(1-\frac{3d(c)}{2}+O(c))\Phi_{t}^{\prime\prime}(x_{j}(t))$
$=x_{j}(t)d(c)(1-3c^{1/2}+O(c))\Phi_{t}^{\prime\prime}(x_{j}(t)).$ (4.6)
Substituting (4.5) and (4.6) back into (4.4), we obtain
$(\Phi_{t}\circ f_{\phi_{t}^{j},c})^{\prime\prime}(x_{j}(t))$ $=\frac{d(c)(1+d(c))}{2\sqrt{1-e^{-c}}}(1-3c^{1/2}+O(c))\Phi_{t}^{\prime\prime}(x_{j}(t)).$
More computations give us $\frac{d(c)(1+d(c))}{2\sqrt{1-e^{-c}}}=1+3c^{1/2}+O(c)$, and so
$(\Phi_{t}\circ f_{\phi_{t}^{j},c})^{\prime\prime}(x_{j}(t))$ $=(1+O(c))\Phi_{t}^{\prime\prime}(x_{j}(t)).$
Therefore, while attaching a slit at $j$, $\frac{\mathrm{d}}{\mathrm{d}t}\log\Phi_{t}^{\prime\prime}(x_{j}(t))=\exp(O(1))$. Together with the previous case, this shows $p^{j}_{t}:=\frac{|\Phi_{t}^{\prime\prime}(x_{j}(t))|^{-\eta}}{\sum_{l}|\Phi_{t}^{\prime\prime}(x_{l}(t))|^{-\eta}}$ is Lipschitz in $t$. Similar arguments apply to the Laplacian path model. ∎
@@ 4.2 Grönwall-type bound on the total error
$\frac{\partial}{\partial t}\bar{x}_{j}(t)$ =$-\sum_{l\not=j}\bar{p}^{j}_{t}\bar{x}_{j}(t)\frac{\bar{x}_{j}(t)+\bar{x}_{l}(t)}{\bar{x}_{j}(t)-\bar{x}_{l}(t)}$ (4.7)
@@
$\frac{\partial}{\partial t}x_{j}(t)$ =$\left\{\begin{matrix}0&\text{if $\xi_{t}=x_{j}(t)$,}\\ -x_{j}(t)\frac{x_{j}(t)+x_{l}(t)}{x_{j}(t)-x_{l}(t)}&\text{if $\xi_{t}=x_{l}(t)$ for $l\not=j$}\end{matrix}\right..$ (4.8)
@@
$=(I^{l}_{n\mathbf{c}}-p^{l}_{n\mathbf{c}})\int_{n\mathbf{c}}^{(n+1)\mathbf{c}\wedge t}\bar{\lambda}_{s}^{j,l}\,\mathrm{d}s+O(\mathbf{c}).$
@@
In the proof of Proposition 4.8 we will use some analytic techniques from [1, Section 4.2]. First we require a few further definitions, mostly following the notation of [1].
@@
Note that if $t=n\mathbf{c}$ for some $n\in\mathbb{N}$, then $h_{0}=f_{1}\circ\dots\circ f_{n}$ and $h_{k\mathbf{c}}=f_{k+1}\circ\dots\circ f_{n}$ for $k\in\mathbb{N}$ with $k<n$.
@@
Let $j\in\{1,\dots,k\}$ and $n\leq\left\lfloor T/\mathbf{c}\right\rfloor$. Let $p>0$ be such that $p^{j}_{m\mathbf{c}}\geq p$ for all $m\leq\left\lfloor T/\mathbf{c}\right\rfloor$ almost surely (we know that such a deterministic constant $p$ must exist). Define the event $H_{j}(n)$ to hold if and only if both $I^{j}_{n\mathbf{c}}=1$ (i.e. $\theta_{n}=\phi_{(n-1)\mathbf{c}}^{j}$) and, for all $m\leq n$,
$\sum_{l=m}^{n}I^{j}_{l\mathbf{c}}\geq\frac{1}{2}p(n-m+1).$
@@
$+\left|\frac{\Phi_{0}^{\prime}(\bar{w}_{j}(0))}{\bar{w}_{j}(0)-\bar{x}_{j}(0)}\right|\left|\kappa_{j}(0)(w_{j}(0)-x_{j}(0))-\bar{\kappa}_{j}(0)(\bar{w}_{j}(0)-\bar{x}_{j}(0))\right|.$ (4.19)
We claim that both $\left|\frac{\Phi_{0}^{\prime}(\bar{w}_{j}(0))}{\bar{w}_{j}(0)-\bar{x}_{j}(0)}\right|$ and $|\kappa_{j}(0)(w_{j}(0)-x_{j}(0))|$ are bounded. The former is bounded by $\sup_{\zeta\in[x_{j}(0),\bar{w}_{j}(0)]}|\Phi_{0}^{\prime\prime}(\zeta)|$. The Schwarz–Christoffel representation [11, equation (1)] tells us $\Phi_{0}^{\prime}(z)=\frac{\Phi_{0}(z)}{z}S(z)$, where $S(z):=\frac{\prod_{l=1}^{k}(z-x_{j}(0))}{\sqrt{\prod_{l=1}^{k}(z-b_{+}^{l})(z-b_{-}^{l})}}$, and $b_{\pm}^{j}$ are the two preimages of the base of the $j$th slit. We can differentiate this to verify that $|\Phi_{0}^{\prime\prime}(\zeta)|$ is bounded when $\zeta$ is not close to any of the $b_{\pm}^{l}$, which is the case for $\zeta\in[x_{j}(0),\bar{w}_{j}(0)]$. Next note that $\sum_{l=1}^{k}|Q_{s}^{j,l}|=O((t-s)^{-1/2})$, which is integrable, and so we can establish a finite upper bound and positive lower bound on $|\kappa_{j}(0)(w_{j}(0)-x_{j}(0))|$ using (4.14) (the proof of which does not rely on (4.13)).
@@
As $t$ is fixed, we will denote $\frac{\partial}{\partial s}h_{s}(z)$ by $\dot{h}_{s}(z)$, etc.. The main effect of conditioning on $H_{j}(n)$ is that $h_{s}^{\prime}(x_{j}(t))=0$ for all $s\in[0,t)$ (the same is always true for the LPM). Where it is unambiguous we will write $x_{j}=x_{j}(s)$, $w_{j}=w_{j}(s)$, etc.. We will prove (4.14), and omit the proof of (4.17) as it is very similar.
@@
$\dot{w}_{j}-\dot{x}_{j}$ =$-\sum_{l}I^{l}_{s}\left(w_{j}+2x_{l}+\frac{2x_{l}^{2}}{w_{j}-x_{l}}\right)+\sum_{l\not=j}I^{l}_{s}x_{j}\frac{x_{j}+x_{l}}{x_{j}-x_{l}}$ (4.22)
=$-I^{j}_{s}\left(w_{j}+2x_{j}+\frac{2x_{j}^{2}}{w_{j}-x_{j}}\right)+\sum_{l\not=j}I^{l}_{s}\left(\frac{2x_{l}^{2}}{(x_{j}-x_{l})(w_{j}-x_{l})}-1\right)(w_{j}-x_{j}),$
@@
where the final inequality comes from the martingale property and the fact that $|I^{l}_{m\mathbf{c}}-p^{l}_{m\mathbf{c}}|\leq 1$ (the factor of 2 is because we bound both the real and imaginary parts of $\bar{Q}^{j,l}_{s}$ by $|\bar{Q}^{j,l}_{s}|$). Then $|\bar{Q}^{j,l}_{s}|\leq A/\sqrt{t-s}$, and so $\sum_{m=0}^{n-1}\mathbb{E}(\int_{m\mathbf{c}}^{(m+1)\mathbf{c}}|\bar{Q}^{j,l}_{s}|\,\mathrm{d}s)^{2}\leq A\mathbf{c}\sum_{m=0}^{n-1}(\sqrt{m+1}-\sqrt{m})^{2}\leq A\mathbf{c}\sum_{m=0}^{n-1}\frac{A}{m+1}\leq A\mathbf{c}\log\frac{t}{\mathbf{c}}$. Define the event
@@
Since the $s$-derivative of $u_{s}$ is singular as $s\uparrow t$, we will carefully analyse and control the singularity. To simplify the notation we will write $w_{j}=w_{j}(s)$, etc.. Subtract (4.22) from its LPM equivalent to obtain
$\dot{u}_{s}=$ $-((w_{j}-x_{j})-(\bar{w}_{j}-\bar{x}_{j}))$ (4.26)
@@
$-I^{j}_{s}\left(3x_{j}+\frac{2x_{j}^{2}}{w_{j}-x_{j}}\right)+\bar{p}^{j}_{s}\left(3\bar{x}_{j}+\frac{2\bar{x}_{j}^{2}}{\bar{w}_{j}-\bar{x}_{j}}\right).$
We will linearise all these terms to write (4.26) in the form $\dot{u}_{s}=\gamma(s)u_{s}+\alpha(s)$ for suitable functions $\gamma$ and $\alpha$. The first line on the right-hand side is clearly $-u_{s}$. The $l$th summand on the second line is a difference of the form $\bar{a}\bar{b}\bar{c}-abc$, so can be linearised as $(\bar{a}-a)\bar{b}\bar{c}+a(\bar{b}-b)\bar{c}+ab(\bar{c}-c)$, giving
@@
If we write $\bar{p}^{l}_{s}=p^{l}_{s}+(\bar{p}^{l}_{s}-p^{l}_{s})$, then the first term becomes the sum of $-(I^{l}_{s}-p^{l}_{s})\frac{2\bar{x}_{l}^{2}(\bar{w}_{j}-\bar{x}_{j})}{(\bar{x}_{j}-\bar{x}_{l})(\bar{w}_{j}-\bar{x}_{l})}$ and something bounded by $A\delta_{\mathrm{total}}(s)$. The final term is also bounded by a multiple of $u_{s}$, and so for two bounded functions $A^{l}_{1}$ and $A^{l}_{2}$ we can write the $l$th summand as
@@
$\dot{u}_{s}$ =$\left(\frac{2I^{j}_{s}x_{j}\bar{x}_{j}}{(w_{j}-x_{j})(\bar{w}_{j}-\bar{x}_{j})}+\sum_{l\not=j}A^{l}_{2}(s)-1\right)u_{s}+A^{j}(s)\frac{\delta_{\mathrm{total}}(s)}{\sqrt{t-s}}$ (4.29)
$\phantom{=}-\sum_{l\not=j}\left((I^{l}_{s}-p^{l}_{s})\frac{2\bar{x}_{l}^{2}(\bar{w}_{j}-\bar{x}_{j})}{(\bar{x}_{j}-\bar{x}_{l})(\bar{w}_{j}-\bar{x}_{l})}\right)-(I^{j}_{s}-p^{j}_{s})\left(3\bar{x}_{j}+\frac{2\bar{x}_{j}^{2}}{\bar{w}_{j}-\bar{x}_{j}}\right).$
@@
$\leq\int_{0}^{t}\frac{|X(s)|}{t-s}\,\mathrm{d}s$ (4.32)
$\phantom{=}+A\underbrace{\int_{0}^{t}\frac{1}{t-s}\int_{s}^{t}\frac{1}{t-r}\int_{r}^{t}|G(v)|\,\mathrm{d}v\,\mathrm{d}r\,\mathrm{d}s}_{(B1)}+A\underbrace{\int_{0}^{t}\frac{1}{t-s}\int_{s}^{t}\frac{|M_{r}|}{t-r}\,\mathrm{d}r\,\mathrm{d}s}_{(B2)}.$

### v3

@@ 5.1 Lipschitz bound on the locations and probabilities
Controlling how $\Phi_{t}^{\prime\prime}(x_{j}(t))$ changes over time will involve a spatial Taylor expansion. Therefore we need to control the third derivative, so first we show that the third derivative at $x_{j}(t)$ takes a particularly convenient form if $\xi_{t-}=x_{j}(t)$. If $f_{s,x}(z)$ attaches a particle of capacity $s>0$ at $x\in\mathbb{T}$, then we can calculate from (3.2) that $f_{s,1}^{\prime\prime}(1)=\frac{1+d(s)}{2\sqrt{1-e^{-s}}}$ and $f_{s,1}^{(3)}(1)=\frac{-3(1+d(s))}{2\sqrt{1-e^{-s}}}=-3f_{s,1}^{\prime\prime}(1)$, therefore $f_{s,x}^{(3)}(x)=-\frac{3}{x}f_{s,x}^{\prime\prime}(x)$. Suppose $\Phi_{t}$ can be written as $\Phi_{t-s}\circ f_{s,x_{j}(t)}$ for some $s\in(0,\mathbf{c}]$, then in this case we have $x_{j}(t-s)=x_{j}(t)$ and $\Phi_{t-s}^{\prime}(x_{j}(t))=0$. Using the chain rule to calculate $(\Phi_{t-s}\circ f_{s,x_{j}(t)})^{\prime\prime}(x_{j}(t))$ and $(\Phi_{t-s}\circ f_{s,x_{j}(t)})^{(3)}(x_{j}(t))$, we also obtain $\Phi_{t}^{(3)}(x_{j}(t))=-\frac{3\Phi_{t}^{\prime\prime}(x_{j}(t))}{x_{j}(t)}$.
Suppose $t,t+\delta\in[m\mathbf{c},(m+1)\mathbf{c}]$ with $\delta>0$. We will show that $|\Phi_{t+\delta}^{\prime\prime}(x_{j}(t+\delta))-\Phi_{t}^{\prime\prime}(x_{j}(t))|=O(\delta)$, which is sufficient to show that $t\mapsto\Phi_{t}^{\prime\prime}(x_{j}(t))$ is Lipschitz on $[0,T]$. There are two cases: either (a) $\xi_{s}=x_{j}(s)$ for all $s\in[m\mathbf{c},(m+1)\mathbf{c})$ or (b) $\xi_{s}=x_{l}(s)$ for some $l\not=j$ over the same interval.
In case (a), $\Phi_{t+\delta}=\Phi_{t}\circ f_{\delta,x_{j}(t)}$. Although we do not necessarily have $\Phi_{t}^{(3)}(x_{j}(t))=-\frac{3\Phi_{t}^{\prime\prime}(x_{j}(t))}{x_{j}(t)}$, we do have $\Phi_{t+\delta^{2}}^{(3)}(x_{j}(t))=-\frac{3\Phi_{t+\delta^{2}}^{\prime\prime}(x_{j}(t))}{x_{j}(t)}$, so let $t^{\prime}=t+\delta^{2}$ and $\delta^{\prime}=\delta-\delta^{2}$, then write $\Phi_{t+\delta}=\Phi_{t^{\prime}}\circ f_{\delta^{\prime},x_{j}(t)}$. To simplify the expressions write $x=x_{j}(t)$. Using the chain rule and the fact that $\Phi_{t^{\prime}}^{\prime}(x)=0$, we obtain
$\Phi_{t+\delta}^{\prime\prime}(x)$ $=\Phi_{t^{\prime}}^{\prime}(f_{\delta^{\prime},x}(x))f_{\delta^{\prime},x}^{\prime\prime}(x)+0$
$=\Phi_{t^{\prime}}^{\prime}(f_{\delta^{\prime},x}(x))\frac{1}{x}\frac{1+d(\delta^{\prime})}{2\sqrt{1-e^{-\delta^{\prime}}}}.$ (5.3)
Since $f_{\delta^{\prime},x}(x)-x=xd(\delta^{\prime})$, by Taylor’s theorem,
$\Phi_{t^{\prime}}^{\prime}(f_{\delta^{\prime},x}(x))$ $=xd(\delta^{\prime})\Phi_{t^{\prime}}^{\prime\prime}(x)+\frac{x^{2}d(\delta^{\prime})^{2}}{2}\Phi_{t^{\prime}}^{(3)}(x)+O(d(\delta^{\prime})^{3})$
$=xd(\delta^{\prime})\left(1-\frac{3}{2}d(\delta^{\prime})\right)\Phi_{t^{\prime}}^{\prime\prime}(x)+O(d(\delta^{\prime})^{3}).$
Then substituting this back into (5.3), we get
$\Phi_{t+\delta}^{\prime\prime}(x)$ $=\frac{d(\delta^{\prime})(1+d(\delta^{\prime}))}{2\sqrt{1-e^{-\delta^{\prime}}}}\left(1-\frac{3}{2}d(\delta^{\prime})\right)\Phi_{t^{\prime}}^{\prime\prime}(x)+O(d(\delta^{\prime})^{2}).$
Since $d(\delta^{\prime})=2e^{\delta^{\prime}}(1+\sqrt{1-e^{-\delta^{\prime}}})-2=2\sqrt{\delta^{\prime}}+O(\delta^{\prime})$, we can calculate $\frac{d(\delta^{\prime})(1+d(\delta^{\prime}))}{2\sqrt{1-e^{-\delta^{\prime}}}}=1+3\sqrt{\delta^{\prime}}+O(\delta^{\prime})$, and $1-\frac{3}{2}d(\delta^{\prime})=1-3\sqrt{\delta^{\prime}}+O(\delta^{\prime})$, so
$\Phi_{t+\delta}^{\prime\prime}(x)$ $=\Phi_{t^{\prime}}^{\prime\prime}(x)+O(\delta^{\prime}).$
Then writing $\Phi_{t^{\prime}}=\Phi_{t}\circ f_{\delta^{2},x}$ and expanding with the chain rule and Taylor’s theorem as above, we do not have as much cancellation and get the slightly cruder expression
$\Phi_{t^{\prime}}^{\prime\prime}(x)=\frac{d(\delta^{2})(1+\delta^{2})}{2\sqrt{1-e^{-\delta^{2}}}}\Phi_{t}^{\prime\prime}(x)+O(d(\delta^{2})),$
but this is good enough. It gives us $\Phi_{t^{\prime}}^{\prime\prime}(x_{j}(t))=\Phi_{t}^{\prime\prime}(x_{j}(t))+O(\delta)$, and so
$\Phi_{t+\delta}^{\prime\prime}(x_{j}(t))-\Phi_{t}^{\prime\prime}(x_{j}(t))=O(\delta).$
In case (b), by the chain rule we have, for $s\in[t,\delta+t)$,
$\frac{\mathrm{d}}{\mathrm{d}s}\Phi_{s}^{\prime\prime}(x_{j}(s))$ $=\left.\frac{\partial}{\partial s}\Phi_{s}^{\prime\prime}(z)\right|_{z=x_{j}(s)}+\Phi_{s}^{(3)}(x_{j}(s))\frac{\mathrm{d}}{\mathrm{d}s}x_{j}(s).$
Differentiating (2.1) twice with respect to $z$, the first term above is
$\left.\frac{\partial}{\partial s}\Phi_{s}^{\prime\prime}(z)\right|_{z=x_{j}(s)}$ $=\Phi_{s}^{(3)}(x_{j}(s))x_{j}(s)\frac{x_{j}(s)+\xi_{s}}{x_{j}(s)-\xi_{s}}+\Phi_{s}^{\prime\prime}(x_{j}(s))\left(1-\frac{2\xi_{s}^{2}}{(x_{j}(s)-\xi_{s})^{2}}\right)$
$=-\Phi_{s}^{(3)}(x_{j}(s))\frac{\mathrm{d}}{\mathrm{d}s}x_{j}(s)+\Phi_{s}^{\prime\prime}(x_{j}(s))\left(1-\frac{2\xi_{s}^{2}}{(x_{j}(s)-\xi_{s})^{2}}\right)$
by (5.2). Therefore
$\frac{\mathrm{d}}{\mathrm{d}s}\Phi_{s}^{\prime\prime}(x_{j}(s))$ $=\Phi_{s}^{\prime\prime}(x_{j}(s))\left(1-\frac{2\xi_{s}^{2}}{(x_{j}(s)-\xi_{s})^{2}}\right),$
which is bounded.
We have now shown that $t\mapsto\Phi_{t}^{\prime\prime}(x_{j}(t))$ is Lipschitz on $[0,T]$, but to conclude that $t\mapsto p^{j}_{t}$ is Lipschitz we need $|\Phi_{t}^{\prime\prime}(x_{j}(t))|$ to be bounded below. This follows from the same argument used in the proof of Proposition 4.7. ∎
@@ 5.2 Grönwall-type bound on the total error
$\frac{\mathrm{d}}{\mathrm{d}t}\bar{x}_{j}(t)$ $=-\sum_{l\not=j}\bar{p}^{j}_{t}\bar{x}_{j}(t)\frac{\bar{x}_{j}(t)+\bar{x}_{l}(t)}{\bar{x}_{j}(t)-\bar{x}_{l}(t)}$ (5.4)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}x_{j}(t)$ $=\left\{\begin{matrix}0&\text{if $\xi_{t}=x_{j}(t)$,}\\ -x_{j}(t)\frac{x_{j}(t)+x_{l}(t)}{x_{j}(t)-x_{l}(t)}&\text{if $\xi_{t}=x_{l}(t)$ for $l\not=j$}\end{matrix}\right.$. (5.5)
@@
$=(I^{l}_{n\mathbf{c}}-p^{l}_{n\mathbf{c}})\int_{n\mathbf{c}}^{(n+1)\mathbf{c}\wedge t}\bar{\lambda}_{s}^{j,l}\,\mathrm{d}s+O(\mathbf{c}^{2}).$
@@
In the proof of Proposition 5.6 we will use some analytic techniques from [2, Section 4.2]. First we require a few further definitions, mostly following the notation of [2].
@@
Several transition maps for an ALE-like cluster are shown in Figure 5. Note that if $t=n\mathbf{c}$ for some $n\in\mathbb{N}$, then $h_{0}=f_{1}\circ\dots\circ f_{n}$ and $h_{k\mathbf{c}}=f_{k+1}\circ\dots\circ f_{n}$ for $k\in\mathbb{N}$ with $k<n$.
@@
[…]
Figure 5: Several transition maps as in Definition 5.2.1 for an ALE-like cluster with ten particles attached to an initial three slits. Notice that none of the transition maps attach the initial three slits to the circle, and that the image of $\mathbb{T}$ under the transition map $h_{t}$ is not monotone in $t$: although the capacity is strictly decreasing, the slits are also moving around.
@@
Let $j\in\{1,\dots,k\}$ and $n\leq\left\lfloor T/\mathbf{c}\right\rfloor$. Let $p>0$ be such that $p^{j}_{m\mathbf{c}}\geq p$ for all $m\leq\left\lfloor T/\mathbf{c}\right\rfloor$ almost surely (we know that such a deterministic constant $p$ must exist). Define the event
$H_{j}(n):=\left\{I^{j}_{n\mathbf{c}}=1\text{ and, for all }m\leq n,\sum_{l=m}^{n}I^{j}_{l\mathbf{c}}\geq\frac{1}{2}p(n-m+1)\right\}.$
Remark.
For the rest of Section 5 we will repeatedly condition on $H_{j}(n)$, so the reader may wish to make a note of Definition 5.2.2 or mark its location.
@@
[…]
We claim that both $\left|\frac{\Phi_{0}^{\prime}(\bar{w}_{j}(0))}{\bar{w}_{j}(0)-\bar{x}_{j}(0)}\right|$ and $|\kappa_{j}(0)(w_{j}(0)-x_{j}(0))|$ are bounded. The former is bounded by $\sup_{\zeta\in[x_{j}(0),\bar{w}_{j}(0)]}|\Phi_{0}^{\prime\prime}(\zeta)|$. The Schwarz–Christoffel representation [6, equation (1)] tells us $\Phi_{0}^{\prime}(z)=\frac{\Phi_{0}(z)}{z}S(z)$, where $S(z):=\frac{\prod_{l=1}^{k}(z-x_{j}(0))}{\sqrt{\prod_{l=1}^{k}(z-b_{+}^{l})(z-b_{-}^{l})}}$, and $b_{\pm}^{j}$ are the two preimages of the base of the $j$th slit. We can differentiate this to verify that $|\Phi_{0}^{\prime\prime}(\zeta)|$ is bounded when $\zeta$ is not close to any of the $b_{\pm}^{l}$, which is the case for $\zeta\in[x_{j}(0),\bar{w}_{j}(0)]$. Next note that $\sum_{l=1}^{k}|Q_{s}^{j,l}|=O((t-s)^{-1/2})$, which is integrable, and so we can establish a finite upper bound and positive lower bound on $|\kappa_{j}(0)(w_{j}(0)-x_{j}(0))|$ using (5.11) (the proof of which does not rely on (5.10)).
@@
As $t$ is fixed, we will denote $\frac{\partial}{\partial s}h_{s}(z)$ by $\dot{h}_{s}(z)$, etc. The main effect of conditioning on $H_{j}(n)$ is that $h_{s}^{\prime}(x_{j}(t))=0$ for all $s\in[0,t)$ (the same is always true for the LPM). Where it is unambiguous we will write $x_{j}=x_{j}(s)$, $w_{j}=w_{j}(s)$, etc. We will prove (5.11), and omit the proof of (5.14) as it is very similar.
@@
$\dot{w}_{j}-\dot{x}_{j}$ $=-\sum_{l}I^{l}_{s}\left(w_{j}+2x_{l}+\frac{2x_{l}^{2}}{w_{j}-x_{l}}\right)+\sum_{l\not=j}I^{l}_{s}x_{j}\frac{x_{j}+x_{l}}{x_{j}-x_{l}}$
$=-I^{j}_{s}\left(w_{j}+2x_{j}+\frac{2x_{j}^{2}}{w_{j}-x_{j}}\right)+\sum_{l\not=j}I^{l}_{s}\left(\frac{2x_{l}^{2}}{(x_{j}-x_{l})(w_{j}-x_{l})}-1\right)(w_{j}-x_{j}),$ (5.19)
@@
where the final inequality comes from the martingale property and the fact that $|I^{l}_{m\mathbf{c}}-p^{l}_{m\mathbf{c}}|\leq 1$ (the factor of 2 is because we bound both the real and imaginary parts of $\bar{Q}^{j,l}_{s}$ by $|\bar{Q}^{j,l}_{s}|$). Then $|\bar{Q}^{j,l}_{s}|\leq A/\sqrt{t-s}$, and so
$\sum_{m=0}^{n-1}\mathbb{E}\left(\int_{m\mathbf{c}}^{(m+1)\mathbf{c}}|\bar{Q}^{j,l}_{s}|\,\mathrm{d}s\right)^{2}\leq A\mathbf{c}\sum_{m=0}^{n-1}(\sqrt{m+1}-\sqrt{m})^{2}\leq A\mathbf{c}\sum_{m=0}^{n-1}\frac{A}{m+1}\leq A\mathbf{c}\log\frac{t}{\mathbf{c}}.$
Define the event
@@
Since the $s$-derivative of $u_{s}$ is singular as $s\uparrow t$, we will carefully analyse and control the singularity. To simplify the notation we will write $w_{j}=w_{j}(s)$, etc. Subtract (5.19) from its LPM equivalent to obtain
$\dot{u}_{s}=$ $-((w_{j}-x_{j})-(\bar{w}_{j}-\bar{x}_{j}))$
@@
$-I^{j}_{s}\left(3x_{j}+\frac{2x_{j}^{2}}{w_{j}-x_{j}}\right)+\bar{p}^{j}_{s}\left(3\bar{x}_{j}+\frac{2\bar{x}_{j}^{2}}{\bar{w}_{j}-\bar{x}_{j}}\right).$ (5.23)
We will linearise all these terms to write (5.23) in the form $\dot{u}_{s}=\gamma(s)u_{s}+\alpha(s)$ for suitable functions $\gamma$ and $\alpha$. The first line on the right-hand side is clearly $u_{s}$. The $l$th summand on the second line is a difference of the form $\bar{a}\bar{b}\bar{c}-abc$, so can be linearised as $(\bar{a}-a)\bar{b}\bar{c}+a(\bar{b}-b)\bar{c}+ab(\bar{c}-c)$, giving
@@
If we write $\bar{p}^{l}_{s}=p^{l}_{s}+(\bar{p}^{l}_{s}-p^{l}_{s})$, then the first term becomes the sum of $(I^{l}_{s}-p^{l}_{s})\frac{2\bar{x}_{l}^{2}(\bar{w}_{j}-\bar{x}_{j})}{(\bar{x}_{j}-\bar{x}_{l})(\bar{w}_{j}-\bar{x}_{l})}$ and something bounded by $A\delta_{\mathrm{total}}(s)$. The final term is also bounded by a multiple of $u_{s}$, and so for two bounded functions $A^{l}_{1}$ and $A^{l}_{2}$ we can write the $l$th summand as
@@
$\dot{u}_{s}$ $=\left(\frac{2I^{j}_{s}x_{j}\bar{x}_{j}}{(w_{j}-x_{j})(\bar{w}_{j}-\bar{x}_{j})}+\sum_{l\not=j}A^{l}_{2}(s)-1\right)u_{s}+A^{j}(s)\frac{\delta_{\mathrm{total}}(s)}{\sqrt{t-s}}$
$\phantom{=}-\sum_{l\not=j}\left((I^{l}_{s}-p^{l}_{s})\frac{2\bar{x}_{l}^{2}(\bar{w}_{j}-\bar{x}_{j})}{(\bar{x}_{j}-\bar{x}_{l})(\bar{w}_{j}-\bar{x}_{l})}\right)-(I^{j}_{s}-p^{j}_{s})\left(3\bar{x}_{j}+\frac{2\bar{x}_{j}^{2}}{\bar{w}_{j}-\bar{x}_{j}}\right).$ (5.26)
@@
$\leq\int_{0}^{t}\frac{|X(s)|}{t-s}\,\mathrm{d}s$
$\phantom{=}+A\underbrace{\int_{0}^{t}\frac{1}{t-s}\int_{s}^{t}\frac{1}{t-r}\int_{r}^{t}|G(v)|\,\mathrm{d}v\,\mathrm{d}r\,\mathrm{d}s}_{(B1)}+A\underbrace{\int_{0}^{t}\frac{1}{t-s}\int_{s}^{t}\frac{|M_{r}|}{t-r}\,\mathrm{d}r\,\mathrm{d}s}_{(B2)}.$ (5.29)